Lavalear Corporation uses a process costing system and traces costs through several processing departments, starting with the Bonding Department. The following is information taken from the Bonding Department’s September production cost report.

Cost Data: Bonding Department
Direct materials costs in beginning inventory, September 1 $ 50,400
Conversion costs in beginning inventory, September 1 36,000
Direct materials costs incurred in September 789,750
Conversion costs incurred in September 787,500
Physical Units: Bonding Department
Units in process, September 1 300
Units transferred out during September 1,500
Units in process, September 30 500
Percentages of Completion: Bonding Department
Direct materials, September 1 40 %
Conversion, September 1 25
Direct materials, September 30 75 %
Conversion, September 30 30

Exercise 18.14 Part a

a. Compute the cost per equivalent unit of direct materials and conversion carried forward from August and assigned to the Bonding Department's beginning inventory on September 1.
Answer #1

Beginning inventory equivalent unit of material = 300*40% = 120

Beginning inventory equivalent unit of conversion = 300*25% = 75

Cost per equivalent unit of material = 50400/120 = 420

Cost per equivalent unit of conversion = 36000/75 = 480